Monica's piggybank contains $8.55 in nickels, dimes, and quarters. The number of dimes is 5 more than 3 times the number of nickels. The number of quarters is 2 times the number of dimes. How many of each kind are there?
Answer by Edwin McCravy(20086)


Monica's piggybank contains $8.55 in nickels, dimes, and quarters. 

.05N + .10D + .25Q = 8.55

Multiply through by 100

5N + 10D + 25Q = 855

Divide through by 5

N + 2D + 5Q = 171

The number of dimes is 5 more than 3 times the number of nickels. 

D = 3N + 5

The number of quarters is 2 times the number of dimes. 

Q = 2D

Solve the system of three equations in three unknowns or variables:

N + 2D + 5Q = 171
D = 3N + 5
Q = 2D

Answer 

 3 nickels  = $ .15
14 dimes    =  1.40
28 quarters =  7.00
-------------------
   total    = $8.55

Edwin